12) Consider the following data on distances traveled by people to visit the local amusement park and calculate the relative frequency for the shortest distance.

Distance

Frequency

1-8 miles

15

9-16 miles

12

17-24 miles

7

25-32 miles

5

33-40 miles

1
Answer #1

Solution,

Distance Frequency Relative Frequency
1 - 8 miles 15 15 / 40 = 0.375 = 37.5%
9 - 16 miles 12 12 / 40 = 0.3 = 30%
17 - 24 miles 7 7 / 40 = 0.175 = 17.5%
25 - 32 miles 5 5 / 40 = 0.125 = 12.5%
33 - 40 miles 1 1 / 40 = 0.025 = 2.5%
Total 40 100%